UT 20150528 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson
Observer AO: N/A
Telescope Operator: SAllanson
LBTO Support: TSargent
Watchdog: RGredel (LUCI-team)

Plan for the night

- This night will again be used to verify LUCI2 readiness for science operations in seeing-limited mode

- Detailed plan depends on observing conditions (cirrus at begin, clear later on),
but will be a mix of mostly short imaging and spectroscopic setups

- DThompson will again be the master performing the observations

Preparations

- APramskiy implemented quite a lot of new features and/or improved old ones
based on the input in particular from DThompson and AConrad

Summary

- executed successfully a wide variety of imaging and spectroscopic programs

- MOS-error once, initialization cleared the problem

- Aladdin-issue (same as during last run in Tucson) => Aladdin was not updated on obs2 with this fix, once fix was installed no further problems

Details

All times are UT
Files are luci2.20150528.NNNN.fits

02:20 Sunset, lots of Cirrus around, should disappear within the next few hours

02:47 Steve will collimate near 12h +39d (NGC 4151)

03:19 Still cloudy, but we can see the guide star and clouds look to be thinning

3:20 ### NGC4151 for LBTB/Pott HKspec 0.5"slit 600s
MOS error: msgID=2575109 2015-05-28

03:20:36 Error in sequence MOS - storage to turnout. WARNING holding current still active at c1m5
Status:
grabber @ -lim, SGs off, hold.curr OFF
rotator @ 341.0 hold.curr OFF
translator @ 130.5 -5436, hold.curr ON
Nothing got grabbed yet, should be able to initialize
Init...OK

03:31 Re-send preset (still cloudy)...guide star not found

Back @ nearby position to remain collimated on brighter star...FWHM~0.6" on guider

03:38 Move mask to FPU so it does not sit in turnout for a long time.

03:51 Re-send preset. In a clearing, but another band of clouds coming...nope!
I can see the guide star, but it is faint (15.74 mag should be OK if clear)
Checked...this is about the best guide star to use.

>>>Bug? I reset the script on the UI and re-committed it, but it did not move the

>>> mask back into turnout. Not sure why.

04:04 ### NGC4151_tell
Guide star found, 0.7" FWHM.
Script did not move the mask into turnout. Restart the scheduler...worked

>>>Bug? Scheduler apparently did not get the information that the mask position had

>>> been changed on the LUCI2 gui.

Re-send...failed (guide star @ bottom of acquisition image, in the dark band)
Steve adjusted the pointing

04:13 Re-send NGC4151_tell, preset successful
  1. sky, 0044 source, 0045 slit, 0046 confirm
  2. 0050 4x10.0s INT G200 HKspec @ 1.93um ABBA
    offset 0 +30 from TelescopeService gui
    reset, skip to spectra and commit
  3. 0054 4x10.0s INT G200 HKspec @ 1.93um ABBA

04:30 ### NGC4151 for LBTB/Pott HKspec 0.5"slit 600s
Preset successful, guide star really faint, resolved?
RTD is stuck downloading an image? Display is white, cannot close windows.
xkill and reopen...same problem
popup says "Failed to update ComboBoxes: Expected keyword not found in header: 002 Plane not ready
xkill and reopen...same problem
xkill and log into luci.luci and reopen there..."start" takes a long time to open display but it worked
RTD only sees the most recent single image!

>>>Bug? RTD has similar problem downloading data as seen in Tucson last run.

We managed to load data by following the path /dataB/readout2/DATA/20150528/ and complete acquisition
  1. 0062 3x50.0s MER/INT spectra of NGC4151
    - Alexander says Aladdin was not updated on obs2 with this fix

05:02 Steve checking the pointing near NGC 4151...OK, so prev guidestar was a galaxy

05:04 ### BS9150_LS0.5_tell
Star is very close to the moon, but guide star is bright and WFS looks OK
  1. sky 0064 src 0065 slit (but won't see it...crossed filters) 66 post offset
    - this did not work, the spillover light in the slit was too faint, star not well centered.

05:19 Re-execute script
  1. sky 0068 src 0069 slit
    Move FW1 OH1060 to clear, take another exposure
  2. new slit image, z band filter only
    Move FW1 back to OH1060
  3. post-offset image looks good
  4. 0073 HKspec AM1.59
  5. 0075 zJspec ...star may be saturated

Aladdin on obs2 updated, switching back to running UI from obs2
It came with the last image displayed

05:37 ### L5Boot1426 14 26 48.42 +35 19 06.72 for LBTO/QSO_SubMM HKspec 1"slit 1hr
  1. sky 0077 src 0078 slit 0079 src+slit
    applied -0.25" X offset by hand, 0080 with this offset savemode was NORMAL! 0080-0085
  2. retake exposure integrated...well centered
  3. script took a post blind-offset exposure.
  4. - 0107 1x300s HKspec on QSO+SubMM galaxy
    Can see the QSO continuum in the differenced spectra

>>>Scriptor update: remove the pause that is automatically inserted after a blind offset.

>>>RTD update: do not send a pause after the alignment offset (longslit or MOS)

07:14 ### HIP71172_tell, does 1"HK and 0.5"HK/zJ spec
  1. src 109 slit (not visible) 110 (z only, manually taken) 111 src+slit
  2. .... 1x60.0s HKspec 1.0"slit saturated

>>>Modification: Attempted to manually reconfigure to check source centration on slit and accidentallymoved in the 1.0"slit. This cost ~12 min for the MOS motion alone. Observers may need to do this sort of operation for sanity checks or problem recovey, but they need a LUCI2-like gui that remains in sync with the current configuration of the instrument. Similar suggestion for a READ2 gui for observers.

I tried editing the XML to remove the preset and change the exposure time to 15 seconds, but
did not change from MER mode (for tellurics it really should be LIR). Skipping this telluric.

07:39 ### J1638+1500 16 38 10.38 +15 00 58.26 for LBTO/z5qso z4.80 H16.6 HKspec 0.5"slit 900s
Guider showing 0.6 arcsec FWHM, allsky cam looks clear
  1. sky 0114 src 0115 slit 0116 src+slit - src there but hard to tell how well centered
  2. 0120 1x240.0s HKspec 0.5"slit
    QSO continuum well-detected across both bands

08:10 ### J1638_tell 16 11 28.74 +16 39 56.40 # V6.09
  1. sky 0122 src 0123 slit (can't see it) 0124 (z only slit) 0125 src+slit
  2. 0127 2x30.0s HKspec saturated
    Modified script to do 20 x 3.0s and it still looks saturated.
  3. 0129 still saturated?
  4. 1x3.0s
>>>Note to self...you cannot pause the head item if some part is labeled "running"

08:38 ### NGC6341_mod.xml - apparently the wrong script...insert pause to abort

08:43 ### NGC6341_mod1.xml (rotator problem corrected) Norris M92
Guider showing 0.7" FWHM, looks clear on allsky cam, 0131 0258
Guider sometimes crept above 1.0 arcsec

Overheads in taking data (from commit to End in the log, so only readout and save involved)
  1. x10.0s took 51sec - 40s data - 2.75s RO = 8.25s overhead others:(53,53,52,53)
  2. x 4.0s took 15sec - 4s data - 2.75s RO = 8.25s overhead others:(17,17,16,17)
  3. x20.0s took 32sec - 20s data - 2.75s RO = 10.25s overhead others:(32,31,31,33)

>>>Request: please provide a breakdown of what is happening during this 8-10 seconds of

overhead on each exposure and assess whether it is reducible or not.

10:51 ### M31rv 00 42 07.99 +40 55 01.1 for LBTO/M31rv EL>25@03:10-->02:50am Imaging
EL~33:49 when I started, rising
Guider showing 1.02
Doing the z band first since Norris/M92 finished a little earlier than predicted
Source visible in un-sky-subtracted exposure in J? (0259)...no

10:57 M31rv_z 0260 0274 15x4.0s LIR/INT z
11:27 M31rv_JHKs 0274 0278 15x4.0s LIR/INT J
0279 0283 15x4.0s LIR/INT H
0284 0288 15x4.0s LIR/INT Ks - file numbers off by one?
First two positions of each filter do not look like they are dithered. Odd.
Looks like the source is visible in H, definitely there in Ks
Ks data are a bit uncollimated because of twilight...lost guide/wfs

>>>Initialize still working on MOS but the status panel is back to green. It should stay "moving"

until the initialization is complete.

Adjusted FM4 a little to move field image to the right ~30 pixels on detector

Twilight flats, N3.75 imaging
...nope, its too late (zJH all saturated or close to it, Ks already @ 15kADU in 2.75s

12:01 Closing dome, insert blind mask, filters, N3.75 camera.

##################################################################

  • Notes for improvement

Filename on readout status panel disappeared when first countdown timer showed up. Stayed off.
- seems to be back working normally now.

Should have an "Abort Script" button if possible (or is it insert pause and abort on GEIRS or READ2 gui?)

"Refresh" button on READ2 gui like as on LUCI2 gui (or is this what "reset" does?)

Need full startup script (opens geirs, UI). Alexander thought it already exists.

"Mask in turnout" message on status panel should be in black (anything but pink on yellow background)

Grating "not stable" message should not be presented the way it currently is. If the grating closes the
loop successfully the panel should remain green. If the grating is OFF target for some period of time
then the panel should turn yellow with a message of "grating lost wavelength stabilization"

The pause the RTD sends to the queue with the alignment offset can be removed

-- %USERSIG{JochenHeidt - 2015-05-28}%
Topic revision: r5 - 11 Jun 2015, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback